# RILe: Reinforced Imitation Learning

### v1

@@ 4 RILe: Reinforced Imitation Learning
At the same time, the teacher agent aims to recover the optimal policy $\pi_{T}^{*}$ as
$\pi_{T}^{*}=\operatorname*{argmax}_{\pi_{T}}E_{\begin{subarray}{c}s^{T}\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\text{log}(D_{\phi}(s^{T}_{t}))\sigma(a^{T})\right].$ (11)
To prove that the student agent can learn expert-like behavior, we need to show that the teacher agent learns to give higher rewards to student experiences that match with the expert state-action pair distribution, as this would enable a student policy to eventually mimic expert behavior.
Lemma 1: Given the discriminator $D_{\phi}$, the teacher agent optimizes its policy $\pi^{\theta_{T}}$ via policy gradients to provide rewards that guide the student agent to match expert’s state-action distributions.
Proof for Lemma 1 The assumptions are presented in Appendix A. The expert’s state-action distribution is denoted by $p_{\text{expert}}(s,a)$. The role of the teacher is to provide a reward signal to the student that encourages the approximation of $p_{\text{expert}}(s,a)$ as closely as possible.
We have $D_{\phi}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ as the discriminator, parameterized by $\phi$, which outputs the likelihood that a given state-action pair $(s,a)$ originates from the expert, as opposed to the student. The teacher’s policy $\pi^{\theta_{T}}$, parameterized by $\theta_{T}$, aims to maximize the likelihood under the discriminator’s assessment, thus encouraging the student agent to generate state-action pairs drawn from a distribution resembling $p_{\text{expert}}(s,a)$.
The Value and Q functions of the teacher, conditioned on the rewards provided by the discriminator, are defined in terms of expected cumulative discriminator rewards. The value function for the teacher’s policy parameters $\theta_{T}$ at state $s_{t}$, simplified from $s^{T}_{t}$ for better readability, is given by:
@@
where $r_{t+k}=D_{\phi}(s_{t+k})$. Similarly, the Q-function for taking action $a_{t}$ in state $s_{t}$ and then following policy $\pi^{\theta_{T}}$ can be written as:
@@
The teacher’s policy optimization is done by maximizing the following clipped surrogate objective function:
@@
where $A^{\theta_{T_{old}}}(s_{t},a_{t})$ is the advantage function computed as $Q^{\theta_{T_{old}}}(s_{t},a_{t})-V^{\theta_{T_{old}}}(s_{t})$, defined with respect to the teacher’s old policy parameters $\theta_{T_{old}}$.
@@
$A^{\theta_{T_{old}}}((s^{S}_{t},a^{S}_{t}),a_{t})=D_{\phi}((s^{S}_{t},a^{S}_{t}))+\gamma\mathbb{E}_{\pi^{\theta_{S}}}\left[V^{\theta_{T_{old}}}((s^{S}_{t+1},a^{S}_{t+1}))\right]-V^{\theta_{T_{old}}}(s^{S}_{t},a^{S}_{t}).$ (16)
@@
The update rule in Equation 17, driven by cumulative rewards from the discriminator, incrementally adapts the teacher’s policy to reinforce student behaviors that are indistinguishable from those of the expert according to $D_{\phi}$. The teacher’s policy is guided to facilitate the student’s approximation of $p_{\text{expert}}(s,a)$, thereby fulfilling its role in the imitation learning process.
@@ 5 Experiments
We evaluate the performance of RILe by addressing three key questions. First, can RILe recover a feasible policy when trained to imitate from noisy expert data? To answer this question, we compare RILe’s performance with common imitation learning baselines in a grid world setting where policies are learned from noisy expert data. Second, can RILe use expert-data explicitly to imitate expert behavior? We use MuJoCo [25, 26] to evaluate RILe’s performance when expert data is leaked to the agents. Third, is RILe efficient and scalable to high-dimensional continuous control tasks? We answer this by using LocoMujoco [27] as the benchmark for evaluating RILe’s effectiveness in imitating motion-capture data within continuous robotic control tasks.
@@
Baselines We compared RILe with four baseline methods: Behavioral cloning (BC [7, 28]), adversarial imitation learning (GAIL [3]), adversarial inverse reinforcement learning (AIRL [4]) and inverse reinforcement learning (IQ-Learn [20]).
5.1 Noisy Expert Data
(a) Expert traj.
(b) RILe traj.
(c) GAIL traj.
(d) AIRL traj.
(e) IQLearn traj.
(f) RILe val.
(g) GAIL val.
(h) AIRL val.
(i) IQLearn val.
Figure 2: In a 5x5 grid environment with lava, (a) the expert trajectory is characterized by noisy data that passes through lava without resulting in death. (c) GAIL, (d) AIRL and (e) IQLearn learn to imitate the expert’s path precisely, leading them to either get stuck near the lava or enter it and perish. (b) RILe avoids the noisy data, better mimics the expert in later stages, and successfully reaches the goal. Subfigures (f-i) display the value tables for RILe, GAIL, AIRL, and IQLearn respectively. The optimal path, derived from the reward of the teacher or discriminator, is highlighted with green lines.
GAIL (Fig. 2c), AIRL (Fig. 2d) and IQLearn (Fig. 2e) fail to reach the goal, as their agents either become stuck or are directed into lava. In contrast, RILe (Fig. 2d) successfully reaches the goal, demonstrating its ability to navigate around imperfections in expert data. The difference in the value graphs between RILe and the baselines intuitively explains this outcome. In AIL and AIRL (Fig. 2f-g), the optimal paths, defined by the actions most rewarded by their discriminators, follow the noisy expert data perfectly. Similarly, in IQLearn, the agent tries to match expert state-actions as closely as possible, minimizing any deviation from the expert trajectory. In contrast, RILe’s teacher agent, trained using RL, adds an extra degree of freedom in the adversarial IL/IRL setting. By providing rewards that maximize cumulative returns from the discriminator, rather than deriving the reward directly from its output, the value graph (Fig. 2f) can learn to circumvent the lava tile in order to follow the expert trajectory to the goal. Consequently, the optimal path of the student agent can overcome the sub-optimal state suggested by the noisy expert demonstration. Since the student agent is guided by the teacher to also match the expert trajectory, it remains close to this path after passing the lava tiles.
5.2 Expert Data Leakage in Training
We investigate whether RILe can use expert data explicitly to imitate expert behavior better. Thus, we analyze how the performance and convergence time will be effected when we leak expert data to the replay buffers of both student and teacher agent. For this experiment, we employed MuJoCo’s Humanoid environment [26, 25] and used a single trajectory from the expert data of [20]. The expert data is leaked to agents’ replay buffers with following percentages: $25\%,50\%,75\%,100\%$. $25\%$ signifies that $25\%$ of the replay buffer’s data is sourced from the expert trajectory, while the remaining $75\%$ originates from the agents’ interactions with the environment. $100\%$ indicates that the agents’ environment interactions are entirely disregarded, with only expert data being used for training.
As shown in Figure 3, leaking expert data to buffers slightly worsened performance but significantly reduced the convergence time. However, when environment interactions were entirely replaced with expert data, the policy’s performance declined substantially. For comparison, results from IQLearn and Behavioral Cloning are also presented, as both methods rely exclusively on expert data for training. Neither method, however, achieved the same level of performance as RILe.
Normalized Scores
Normalized Steps
Figure 3: Scores (left) representing the total rewards achieved, and convergence steps (right) representing the number of steps to convergence, are presented against the percentage of leaked expert data. For RILe, the x-axis shows the percentage of leaked expert data in the student agent’s buffer, while the y-axis shows it in the teacher agent’s buffer. Scores/Steps are normalized by no-leakage training results of RILe. Behavioral Cloning fails to achieve a comparable score to RILe. While IQLearn converges faster than the no-leakage RILe, it still falls short in terms of overall performance.
@@ 5.3 Motion-Capture Data Imitation for Robotic Continuous Control
In LocoMujoco, the objective is to learn policies that can accurately imitate the collected MoCap data of various robotic tasks. LocoMujoco is challenging due its high dimensionality. The MoCap data only consists of states of the robot, not its actions, which prevents the use of Behavioral Cloning and IQLearn in this analysis, since they both requires to have full expert data with states and actions. It should be noted that LocoMujoco environments do not have terminal or goal conditions.
Results are presented in Table 1 for both test seeds and validation seeds. RILe outperforms baselines especially for new initial conditions introduced by varying seeds.
@@ 6 Discussion
We have demonstrated through our experiments that RILe outperforms baseline models across various settings. While RILe surpasses the state-of-the-art adversarial IL/IRL methods, it also has its limitations. The main challenge is training the policy with a changing reward function, which leads to unstable policy updates for the student agent. To address this issue, we formulate the teacher’s reward by correlating its actions with the output of the discriminator, instead of only relying on the discriminator. However, instability remains a challenge and causes slower learning of the action policy in RILe compared to AIL. Future work should establish bounds for the reward agent updates to enhance the stability of the student agent’s learning process.
Moreover, balancing the learning rates of the discriminator and the policies is challenging, particularly because the discriminator tends to overfit the data quickly. This overfitting makes it difficult for the teacher agent to find a reward for the student that can challenge the discriminator. While adjusting the discriminator’s learning rate on a case-by-case basis helps, future research should focus on a fundamental solution, such as replacing the discriminator with a more robust similarity component.
Uniquely in RILe, the discriminator guides the reward learning process instead of directly guiding policy learning. Therefore, the reward function is learned, not directly derived from the discriminator. This distinction allows RILe to provide meaningful guidance even when the discriminator overfits to noisy expert data. In the LocoMujoco benchmark, where perfect expert data is available, we argue that RILe’s ability to account for the agent’s exploration efforts during training, in addition to its behavioral similarity with the expert, results in a superior generalization capability compared to AIL and AIRL.
@@ Appendix A Justification of RILe
  • For the teacher’s and student’s policy functions $(\pi^{\theta_{T}})$ and $(\pi^{\theta_{S}})$, and the Q-functions $(Q^{\theta_{S}})$, each is Lipschitz continuous with respect to its parameters with constants $(L_{\theta_{T}}),(L_{\theta_{S}}),and(L_{Q})$, respectively. This means for all $(s,a)$ and for any pair of parameter settings $(\theta,\theta^{\prime}):[|\pi^{\theta}(s,a)-\pi^{\theta^{\prime}}(s,a)|\leq L_{\theta}|\theta-\theta^{\prime}|,][|Q^{\theta}(s,a)-Q^{\theta^{\prime}}(s,a)|\leq L_{Q}|\theta-\theta^{\prime}|.]$
@@ A.1 Lemma 2:
Appendix B Compute Resources

### v2

@@ 4 RILe: Reinforced Imitation Learning
At the same time, the trainer agent aims to recover the optimal policy $\pi_{T}^{*}$ as
$\pi_{T}^{*}=\operatorname*{arg\,max}_{\pi_{T}}E_{\begin{subarray}{c}s^{T}\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\upsilon(D_{\phi}(s^{T}_{t}))a^{T}\right].$ (10)
To prove that the student agent can learn expert-like behavior, we need to show that the trainer agent learns to give higher rewards to student experiences that match with the expert state-action pair distribution, as this would enable a student policy to eventually mimic expert behavior.
Lemma 1: Given the discriminator $D_{\phi}$, the trainer agent optimizes its policy $\pi^{\theta_{T}}$ via policy gradients to provide rewards that guide the student agent to match expert’s state-action distributions.
Proof for Lemma 1 The assumptions are presented in Appendix A. The expert’s state-action distribution is denoted by $p_{\text{expert}}(s,a)$. The role of the trainer is to provide a reward signal to the student that encourages the approximation of $p_{\text{expert}}(s,a)$ as closely as possible.
We have $D_{\phi}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ as the discriminator, parameterized by $\phi$, which outputs the likelihood that a given state-action pair $(s,a)$ originates from the expert, as opposed to the student. The trainer’s policy $\pi^{\theta_{T}}$, parameterized by $\theta_{T}$, aims to maximize the likelihood under the discriminator’s assessment, thus encouraging the student agent to generate state-action pairs drawn from a distribution resembling $p_{\text{expert}}(s,a)$.
The Value and Q functions of the trainer, conditioned on the rewards provided by the discriminator, are defined in terms of expected cumulative discriminator rewards. The value function for the trainer’s policy parameters $\theta_{T}$ at state $s_{t}$, simplified from $s^{T}_{t}$ for better readability, is given by:
@@
where $r_{t+k}=\upsilon(D_{\phi}(s_{t+k}))a_{t+k}$. Similarly, the Q-function for taking action $a_{t}$ in state $s_{t}$ and then following policy $\pi^{\theta_{T}}$ can be written as:
@@
The trainer’s policy optimization is done by maximizing the following clipped surrogate objective function:
@@
where $A^{\theta_{T_{old}}}(s_{t},a_{t})$ is the advantage function computed as $Q^{\theta_{T_{old}}}(s_{t},a_{t})-V^{\theta_{T_{old}}}(s_{t})$, defined with respect to the trainer’s old policy parameters $\theta_{T_{old}}$.
@@
The update rule in Equation 15, driven by cumulative rewards from the discriminator, incrementally adapts the trainer’s policy to reinforce student behaviors that are indistinguishable from those of the expert according to $D_{\phi}$. The trainer’s policy is guided to facilitate the student’s approximation of $p_{\text{expert}}(s,a)$, thereby fulfilling its role in the imitation learning process.
@@ 5 Experiments
We evaluate the performance of RILe by addressing three key questions:
  1. 1.
    How does RILe’s adaptive reward function evolve compared to baseline methods and how does this evolution enhance the learning process?
  2. 2.
    Is RILe efficient and scalable to high-dimensional continuous control tasks?
  3. 3.
    Can RILe use expert-data explicitly to imitate expert behavior?
To answer the first question, we compare RILe’s performance with AI(R)L baselines in a maze setting, where we demonstrate how the trainer agent modifies the reward function to guide the student during training. For the second question, we evaluate RILe’s effectiveness in imitating motion-capture data within continuous robotic control tasks, using LocoMujoco [20] as the environment. To answer the last question, we use a simulated humanoid character from MuJoCo [21, 22] to evaluate RILe’s performance when expert data is explicitly used by the agents. Additional experimental results are provided in the Appendix, where we evaluate the robustness of the learned reward function in changing environments and analyze the noise resilience of our method.
@@
Baselines We compare RILe with four baseline methods: Behavioral cloning (BC [7, 23]), adversarial imitation learning (GAIL [3]), adversarial inverse reinforcement learning (AIRL [6]) and inverse reinforcement learning (IQ-Learn [24]).
5.1 Evolving Reward Function
To evaluate the impact of RILe’s trainer agent on the learning process compared to existing methods, we designed a maze experiment. The task is to navigate from a fixed start on the left side to a goal in the top-right corner, avoiding obstacles.
Using a single expert demonstration, we trained RILe, GAIL, and AIRL. Fig. 3 shows how the student’s reward function evolves during training with all three methods. For RILe, we plot the reward function learned by the trainer. For GAIL, we visualize the discriminator output, and for AIRL, the reward term under the discriminator. The columns represent reward plots at 25%, 50%, 75%, and 100% of training completion for each algorithm, respectively. We present the student agent’s trajectory from the previous epoch on reward plots to demonstrate how reward functions adapt to the student agent’s progress.
(a) RILe
(b) GAIL
(c) AIRL
Figure 3: Evolution of reward functions during training for (a) RILe, (b) GAIL, and (c) AIRL in a continuous maze environment. Columns show reward landscapes at 25%, 50%, 75%, and 100% of training completion (left to right). The expert’s trajectory is shown in red, while the student agent’s trajectory from the previous training epoch is in black. Color gradients represent reward values, with darker colors indicating lower rewards and brighter colors indicating higher rewards. Purple squares represent obstacles. RILe demonstrates a more adaptive reward function that evolves with the student’s progress, while GAIL and AIRL maintain relatively static reward landscapes throughout training.
The discriminator-based reward functions of GAIL and AIRL remain relatively static, largely independent of the student agent’s current performance. This static behavior is because the discriminators in GAIL and AIRL focus on distinguishing expert data and progressively improve throughout training. In contrast, the trainer in RILe adapts itself based on the current learning progress of the student’s policy.
As the first column shows, when the student’s trajectory does not resemble the expert, RILe’s trainer encourage exploration towards the expert path. This indicates that the trainer can provide informative rewards despite potentially low discriminator feedback. Then, in the second column, as the student progresses, RILe’s reward function adapts dynamically. When the student learns to reach the bottom-right, the trainer shifts high rewards to the top-left, guiding the agent to explore that area. Third column shows as the student approaches the goal, the trainer increases rewards around it while maintaining rewards in specific areas (e.g., the left part) to prevent the agent from getting stuck. Finally, as the fourth column shows, the trainer focuses high rewards on the top-right area near the goal while maintaining a path of rewards that guides the student along the expert trajectory.
RILe’s evolving reward function demonstrates its ability to provide meaningful guidance even when the discriminator easily identifies non-expert policies. By adapting to the student’s current capabilities, RILe creates a dynamic learning curriculum that encourages exploration and gradual improvement towards expert-like behavior. Our results demonstrate that the trainer effectively associates rewards with future behavior improvements, even if delayed. It provides incremental rewards for actions moving the student in the right direction, and adapts the reward function to match the student’s abilities, gradually guiding them toward expert behavior.
@@ 5.2 Motion-Capture Data Imitation for Robotic Continuous Control
We evaluate RILe’s performance on the LocoMujoco benchmark, which presents a challenging task of imitating motion-capture (MoCap) data for various robotic control tasks. This benchmark is particularly demanding due to its high dimensionality and the absence of action data in the MoCap recordings which precludes the use of methods like Behavioral Cloning and IQLearn that require complete state-action pairs.
@@
Table 1 presents the results for seven LocoMujoco tasks, comparing RILe with GAIL and AIRL across both test and validation seeds. The expert performance, set at 1000 for all tasks, serves as the upper benchmark. RILe demonstrates superior performance in all scenarios, particularly excelling in generalization to new initial conditions as evidenced by the test seed results.
5.3 Impact of Expert Data on Trainer-Student Dynamics
Figure 4: Impact of explicit usage of expert data on RILe’s performance. Red markers and yellow markers show normalized scores and normalized steps, respectively. Square markers show IQ-Learn and BC results for comparisons. The percentage of expert data in the buffers are presented around the figure. More expert data in buffers speeds convergence but can reduce final performance, especially at 100%, showing how the explicit usage of excess expert data hinders trainer-student adaptation in RILe
We study how explicitly incorporating expert data into RILe’s training affects the trainer’s ability to adapt to the student’s needs, analyzing performance and convergence time when expert data is added to the agents’ replay buffers. For this experiment, we used MuJoCo’s Humanoid environment [22, 21] and a single expert trajectory from [24]. We varied the proportion of expert data in the replay buffers from 0% to 100%; for example, 25% means a quarter of the buffer is expert data and 75% is from agent interactions.
As shown in Fig. 4, introducing expert data led to faster convergence but decreased performance. Notably, when environmental interactions were completely replaced by the expert data (100% case), the student’s performance declined significantly. Excessive expert data appears to hinder the trainer’s ability to adapt to the student, disrupting RILe’s dynamic learning process. For comparison, we include results from IQ-Learn and Behavioral Cloning (BC), which rely explicitly on expert data. Despite their exclusive use of expert data, neither method matched RILe’s performance, even when RILe used substantial amounts of expert data.
Our findings underscore the importance of balancing environmental interactions and expert demonstrations in buffers in RILe. While some expert data accelerates learning, too much impedes the trainer’s adaptability, underscoring the unique strength of RILe’s interactive learning approach.
@@ 6 Discussion
Our experiments have demonstrated that RILe consistently outperforms baseline models across various settings, surpassing state-of-the-art adversarial IL/IRL methods. RILe’s key strength lies in its adaptive learning approach, where the trainer agent dynamically adjusts the reward function based on the student’s current capabilities. This enables more efficient learning, particularly in complex environments.
However, RILe faces challenges, primarily in maintaining policy stability with a changing reward function. Our solution of freezing the trainer after the student achieves half the expert score is effective but potentially limiting. Additionally, the discriminator’s tendency to overfit quickly presents another significant challenge. Future work should focus on establishing bounds for reward agent updates to enhance stability without compromising adaptability, developing a stronger theoretical understanding of RILe’s convergence properties, and exploring discriminator-less approaches.
Despite these challenges, RILe has demonstrated advantages in adaptability, robustness, and generalization, thereby making it a promising direction for future research in imitation learning and opening up new possibilities for dynamic and responsive learning processes.
@@ Appendix A Justification of RILe
  • For the trainer’s and student’s policy functions $(\pi^{\theta_{T}})$ and $(\pi^{\theta_{S}})$, and the Q-functions $(Q^{\theta_{S}})$, each is Lipschitz continuous with respect to its parameters with constants $(L_{\theta_{T}}),(L_{\theta_{S}}),and(L_{Q})$, respectively. This means for all $(s,a)$ and for any pair of parameter settings $(\theta,\theta^{\prime}):[|\pi^{\theta}(s,a)-\pi^{\theta^{\prime}}(s,a)|\leq L_{\theta}|\theta-\theta^{\prime}|,][|Q^{\theta}(s,a)-Q^{\theta^{\prime}}(s,a)|\leq L_{Q}|\theta-\theta^{\prime}|.]$
@@ A.1 Lemma 2:
Appendix B Additional Experiments
B.1 Noisy Expert Data
(a) Expert traj.
(b) RILe traj.
(c) GAIL traj.
(d) AIRL traj.
(e) IQLearn traj.
(f) RILe val.
(g) GAIL val.
(h) AIRL val.
(i) IQLearn val.
Figure 5: In a 5x5 grid environment with lava, (a) the expert trajectory is characterized by noisy data that passes through lava without resulting in death. (c) GAIL, (d) AIRL and (e) IQLearn learn to imitate the expert’s path precisely, leading them to either get stuck near the lava or enter it and perish. (b) RILe avoids the noisy data, better mimics the expert in later stages, and successfully reaches the goal. Subfigures (f-i) display the value tables for RILe, GAIL, AIRL, and IQLearn respectively. The optimal path, derived from the reward of the trainer or discriminator, is highlighted with green lines.
GAIL (Fig. 5c), AIRL (Fig. 5d) and IQLearn (Fig. 5e) fail to reach the goal, as their agents either become stuck or are directed into lava.
In contrast, RILe (Fig. 5d) successfully reaches the goal, demonstrating its ability to navigate around imperfections in expert data. The difference in the value graphs between RILe and the baselines intuitively explains this outcome. In AIL and AIRL (Fig. 5f-g), the optimal paths, defined by the actions most rewarded by their discriminators, follow the noisy expert data perfectly. Similarly, in IQLearn, the agent tries to match expert state-actions as closely as possible, minimizing any deviation from the expert trajectory. In contrast, RILe’s trainer agent, trained using RL, adds an extra degree of freedom in the adversarial IL/IRL setting. By providing rewards that maximize cumulative returns from the discriminator, rather than deriving the reward directly from its output, the value graph (Fig. 5f) can learn to circumvent the lava tile in order to follow the expert trajectory to the goal. Consequently, the optimal path of the student agent can overcome the sub-optimal state suggested by the noisy expert demonstration. Since the student agent is guided by the trainer to also match the expert trajectory, it remains close to this path after passing the lava tiles.
B.2 Robustness to Noise in the Expert Data
To evaluate the robustness of RILe and baseline methods to noise in the expert data, we conducted experiments in the MuJoCo Humanoid-v2 environment. Artificial noise sampled from a zero-mean Gaussian distribution with varying standard deviations ($\Sigma$) was added to a single expert trajectory, affecting either the actions or the states. The baselines used for comparison were GAIL [3], AIRL [6], RIL-Co [25], IC-GAIL [26], and IQ-Learn [24].
As shown in Table 2, RILe consistently outperforms the baselines across different noise levels, demonstrating superior robustness even when a high amount of noise is present in the expert data ($\Sigma=0.5$). These results indicate that RILe is less sensitive to imperfections in the expert demonstrations compared to existing methods.
Table 2: Test results in MuJoCo Humanoid-v2 environment, where artificial noise sampled from a zero-mean Gaussian distribution is added to a single expert trajectory. Results are aggregated over 20 different-seed environments. IQ-Learn* is trained using the official code and hyperparameters of the IQ-Learn algorithm.
Noise-Free Action Noise State Noise
$\Sigma=0$ $\Sigma=0.2$ $\Sigma=0.5$ $\Sigma=0.2$ $\Sigma=0.5$
RILe 5681 5280 5154 5350 5205
GAIL 5430 5275 902 5147 917
AIRL 5276 4869 4589 4898 4780
RIL-Co 576 491 493 505 501
IC-GAIL 610 601 568 590 591
IQ-Learn* 312 192 153 243 277
B.3 Robustness of the Learned Reward Function
We evaluated the robustness of the reward functions learned by RILe and AIRL [6] through an experiment similar to that conducted by [27]. Initially, both methods were trained to learn reward functions in a noise-free MuJoCo Humanoid-v2 environment. After training, these reward functions were frozen. Subsequently, new student agents were trained using these fixed reward functions in environments where Gaussian noise was added to the agents’ actions, with varying noise levels.
Table 3 presents the results of this evaluation. The reward function learned by RILe demonstrates superior robustness to noise, maintaining high performance even under increased noise levels. In contrast, the performance of agents using the reward function learned by AIRL decreases more significantly as noise increases. These findings indicate that the reward function learned by RILe is more resilient to environmental noise, contributing to better agent performance in noisy conditions.
Table 3: We test the robustness of learned reward functions. After training reward functions in a noise-free setting, reward functions are frozen, and used to train a new agent in a noisy environment, where Gaussian noise is added to agent’s actions in every step.
No Noise Mild Noise High Noise
$\Sigma=0$ $\Sigma=0.2$ $\Sigma=0.5$
RILe 5748 5201 5196
AIRL 5334 5005 4967
B.4 Reward Curves
We compare the reward curves of RILe, GAIL [3], AIRL [6], IQ-Learn [24], and AdapMen [28] in the MuJoCo Humanoid-v2 experiment. Since the task involves learning from expert trajectories, we combined AdapMen with an adversarial discriminator to enable training without an expert policy.
As shown in the reward curves, despite RILe having multiple components, it is the most efficient method. This efficiency is achieved through the dynamic guidance of the trainer during training, which adapts the reward function to meet the student’s needs.
Figure 6: Training reward curves for the MuJoCo Humanoid-v2 experiment comparing RILe, AIRL, GAIL, IQ-Learn*, and adapted AdapMen. AdapMen is combined with an adversarial discriminator to be able to train it without expert policy.
Appendix C Compute Resources